# A variational approach to dimension-free self-normalized concentration

### v1

@@ Appendix B Omitted Proofs
$h_{\tau}(\lambda^{*}(\eta^{j(\tau)+1}))=\frac{\|U_{0}\|_{{\textnormal{op}}}^{1/2}}{g_{\psi,\tau}(\lambda^{*}(\eta^{j(\tau)+1}))}\lesssim\frac{1}{\sqrt{1-1/\gamma_{\min}(V_{\tau})}}.$
@@
$\|S_{\tau}\|_{V_{\tau}^{-1}}$ $\lesssim(\psi^{*})^{-1}\left(\frac{1}{\sqrt{1-1/\gamma_{\min}(V_{\tau})}}\left\{\log(\det V_{\tau})+\log\bigg{(}\frac{\ell(j(\tau))}{\delta}\bigg{)}\right\}\right)$
$\lesssim(\psi^{*})^{-1}\left(\frac{1}{\sqrt{1-1/\gamma_{\min}(V_{\tau})}}(\log(\det V_{\tau})+\log(\log_{\eta}(\|S_{\tau}\|_{V_{\tau}^{-1}})/\delta))\right),$
where we’ve used that $\ell(j(\tau))\asymp j(\tau)^{2}$ and $\log(j(\tau)^{2})\asymp\log(\log_{\eta}\|S_{\tau}\|_{V_{\tau}^{-1}})$, which completes the proof.
B.5 Proof of Corollary 4.8
Since $\psi$ is CGF-like, so too is $\psi^{*}$ [60, Proposition A.1]. In particular, $\psi^{*}$ is strictly convex and increasing, implying that $\varphi$ is strictly increasing and concave. It follows that for all $x,y$,
@@
(A concave function is bounded by its first order Taylor approximation.) Set
$A_{\tau}=\frac{\log(\det V_{\tau})+\log(1/\delta)}{\sqrt{1-1/\gamma_{\min}(V_{\tau})}},\quad B_{\tau}=\frac{\log\log(\|S_{\tau}\|_{V_{\tau}^{-1}})}{\sqrt{1-1/\gamma_{\min}(V_{\tau})}}.$
Since $1-1/\gamma_{\min}(V_{\tau})\leq 1$ we can bound $A_{\tau}$ as $A_{\tau}\geq A_{\min}:=\log(1/\delta)$. Further, since $V_{\tau}\succeq U_{0}$, $\gamma_{\min}(V_{\tau})\geq\gamma_{\min}(U_{0}):=\rho$, hence
$B_{\tau}\leq\frac{\log\log(\|S_{\tau}\|_{V_{\tau}^{-1}})}{\sqrt{1-1/\rho}}\leq\frac{\|S_{\tau}\|_{V_{\tau}^{-1}}}{\sqrt{1-1/\rho}}.$
Take $x=A_{\tau}$ and $y=B_{\tau}$ in (B.31), we obtain that the right hand side of (4.9) is bounded as
$\varphi(A_{\tau}+B_{\tau})$ $\leq\varphi(A_{\tau})+\varphi^{\prime}(A_{\tau})B_{\tau}=\varphi(A_{\tau})+\frac{B_{\tau}}{(\psi^{*})^{\prime}(\varphi(A_{\tau}))}.$
Since $\psi^{*}$ is strictly increasing and convex, $(\psi^{*})^{\prime}$ is also increasing. Therefore the term $(\psi^{*})^{\prime}(\varphi(x))$ is increasing in $x$, and
$\varphi^{\prime}(A_{\tau})\leq\frac{1}{(\psi^{*})^{\prime}(\varphi(A_{\tau}))}\leq\frac{1}{(\psi^{*})^{\prime}(\varphi(\log(1/\delta)))}=\varphi^{\prime}(\log(1/\delta)),$
whence
$\varphi(A_{\tau}+B_{\tau})\leq\varphi(A_{\tau})+\frac{\varphi^{\prime}(\log(1/\delta))\|S_{\tau}\|_{V_{\tau}^{-1}}}{\sqrt{1-1/\rho}}.$
@@ Appendix C Simulation Details
We take $U_{0}=I_{d}$ and $d=20$. For the first 100 time steps, we make a full rank update in order t o grow the minimum eigenvalue sufficiently. These updates are dampened based on the size of the eigenvalue, however, so there is still anisotropic growth of the matrix. The dampening factor is 1 for the top eigenvector (i.e., corresponding to the largest eigenvalue) and 0.1 for the bottom, with a linear interpolation between 1 and 0.1 in the remaining directions.
@@
where $u_{1}$ is the eigenvector corresponding to the largest eigenvalue, $u_{2}$ that corresponding to the second largest eigenvalue, and so on. For small growth of the determinant (Figure 2(a)) we use $k=1$, for moderate growth (Figure 2(b)) we use $k=4$ and for large growth (Figure 2(c)) we use $k=8$.
We compare Theorem 4.3 to Corollary 2 in Whitehouse et al. [60], which is tailored explicitly to sub-$\psi_{G,c}$ processes. It states that if $(S_{t},V_{t})$ is a sub-$\psi_{G,c}$ process with $V_{t}\succeq\rho I_{d}$ for all $t$, then, with probability $1-\delta$, for all $t$,
@@
$M_{1}(t)$ $=B\log\left(A\log_{\eta}\left(\frac{\gamma_{\max}(V_{t})}{\rho}\right)\right),$
@@
We take $\beta=2$, $\eta=2$, and $\epsilon=1/2$, as suggested in their paper. In all plots, we omit the first 200 time steps because our bound blows up when $\gamma_{\min}(V_{t})$ is too small. As we say in the main paper, this is drawback of our bound from which the bound of Whitehouse et al. [60] does not suffer.
@@
For Figure 3 we generate $V_{t}$ in the same way as in Figure 2. We use $k=5$, $d=20$, and $U_{0}=I_{d}$. We use $c=1/4$ as the sub-gamma parameter. Note that $\psi_{G,1/4}^{-1}(1)\approx 1.54$ so our choices of $\lambda$ are all legal.

### v2

@@ Appendix B Omitted Proofs
$h_{\tau}(\lambda^{*}(\eta^{j(\tau)+1}))\lesssim\frac{1}{\sqrt{1-\gamma_{\max}(U_{0})/\gamma_{\min}(\widehat{V}_{\tau})}}.$
@@
$\|S_{\tau}\|_{V_{\tau}^{-1}}$ $\lesssim(\psi^{*})^{-1}\left(\frac{\log(\det V_{\tau})+\log\bigg(\frac{\ell(j(\tau))}{\delta}\bigg)}{\sqrt{1-\gamma_{\max}(U_{0})/\gamma_{\min}(\widehat{V}_{\tau})}}+f(2\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})\right)$
$\lesssim(\psi^{*})^{-1}\left(\frac{\log(\det V_{\tau})+\log(\log_{\eta}(\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})/\delta)}{\sqrt{1-\gamma_{\max}(U_{0})/\gamma_{\min}(\widehat{V}_{\tau})}}+f(2\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})\right),$
where we’ve used that $\ell(j(\tau))\asymp j(\tau)^{2}$ and $\log(j(\tau)^{2})\asymp\log(\log_{\eta}\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})$, which completes the proof.
B.6 Proof of Corollary 4.10
Since $\psi$ is CGF-like, so too is $\psi^{*}$ [63, Proposition A.1]. In particular, $\psi^{*}$ is strictly convex and increasing, implying that $\varphi$ is strictly increasing and concave. It follows that for all $x,y$,
@@
(A concave function is bounded by its first order Taylor approximation.) Let $\widehat{V}_{t}=V_{t}+U_{0}$ and set
$A_{\tau}=\frac{\log(\det\widehat{V}_{\tau})+\log(1/\delta)}{\sqrt{1-\upsilon/\gamma_{\min}(\widehat{V}_{\tau})}},\quad B_{\tau}=\frac{\log\log(\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})}{\sqrt{1-\upsilon/\gamma_{\min}(\widehat{V}_{\tau})}}+f(2\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}).$
Now, for $\tau>\tau_{0}=\inf\{t:\upsilon/\gamma_{\min}(\widehat{V}_{t})\leq b\}$ we have $1-\upsilon/\gamma_{\min}(\widehat{V}_{\tau})\geq 1-b$. For such $\tau$ therefore, we have
$B_{\tau}\leq\frac{\log\log(\|S_{\tau}\|_{V_{\tau}^{-1}})}{\sqrt{1-1/\upsilon}}+f(2\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})\leq\frac{\|S_{\tau}\|_{V_{\tau}^{-1}}(1+2\kappa)}{\sqrt{1-b}},$
using the assumption that $0<b<1$. Taking $x=A_{\tau}$ and $y=B_{\tau}$ in (B.32), we obtain that the right hand side of (4.13) is bounded as
$\varphi(A_{\tau}+B_{\tau})$ $\leq\varphi(A_{\tau})+\varphi^{\prime}(A_{\tau})B_{\tau},$
where we’ve used that $A_{\tau}\geq\log(1/\delta)$ for $\tau\geq\tau_{0}$. Since $\psi^{*}$ is strictly increasing and convex, $(\psi^{*})^{\prime}$ is also increasing. Therefore the term $(\psi^{*})^{\prime}(\varphi(x))$ is increasing in $x$, and
$\varphi^{\prime}(A_{\tau})=\frac{1}{(\psi^{*})^{\prime}(\varphi(A_{\tau}))}\leq\frac{1}{(\psi^{*})^{\prime}(\varphi(\log(1/\delta)))}=\varphi^{\prime}(\log(1/\delta)),$
where the equalities follows from the chain rule. Hence, with probability $1-\delta$, whenever $\tau\geq\tau_{0}$,
$\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}\leq\varphi(A_{\tau}+B_{\tau})\leq\varphi(A_{\tau})+\frac{\varphi^{\prime}(\log(1/\delta))\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}(1+2\kappa)}{\sqrt{1-b}}.$
@@ Appendix C Simulation Details
We take $U_{0}=I_{d}$ and $d=20$ and $c=1$. For the first 100 time steps, we make a full rank update in order to sufficiently grow the minimum eigenvalue. These updates are dampened based on the size of the eigenvalue, however, so there is still anisotropic growth of the matrix. The dampening factor is 1 for the top eigenvector (i.e., corresponding to the largest eigenvalue) and 0.1 for the bottom, with a linear interpolation between 1 and 0.1 in the remaining directions.
@@
where $u_{1}$ is the eigenvector corresponding to the largest eigenvalue, $u_{2}$ that corresponding to the second largest eigenvalue, and so on. For small growth of the determinant (the left panel) we use $k=1$, for moderate growth (middle panel) we use $k=10$ and for large growth (right panel) we use k=19=d1k=19=d-1 (nearly a full-rank update).
We compare Theorem 4.4 to Corollary 2 in Whitehouse et al. [63], which is tailored explicitly to sub-$\psi_{G,c}$ processes. It states that if $(S_{t},V_{t})$ is a sub-$\psi_{G,c}$ process with $V_{t}\succeq\upsilon I_{d}$ for all $t$, then, with probability $1-\delta$, for all $t$,
@@
$M_{1}(t)$ $=B\log\left(A\log_{\eta}\left(\frac{\gamma_{\max}(V_{t})}{\upsilon}\right)\right),$
@@
We take $\beta=2$, $\eta=2$, and $\epsilon=1/2$, as suggested in their paper. In all plots, we omit the first 100 time steps because our bound blows up when $\gamma_{\min}(V_{t})$ is too small. As we say in the main paper, this is drawback of our bound, and one from which the bound of Whitehouse et al. [63] does not suffer. Note that $\psi_{G,1}^{-1}(1)\approx 0.73$, so all our choices of $\lambda$ are legal.
@@
We take $d=20$ and $U_{0}=2\cdot I_{d}$. For $k\in\{2,5,10\}$, we generate $X_{t}$ by letting the first $k$ components be random uniform noise, and then normalizing the vector such that $\|X_{t}\|_{2}\leq 1/2$. We take $\widehat{\mu}_{t}$ to be the empirical mean of the first $t$ observations, i.e., $\widehat{\mu}_{t}=t^{-1}\sum_{k\leq t}X_{k}$.
A per the discussion after Corollary 6.1, when instantiating our bound we split $U_{0}$ into $\Gamma$ and $U_{0}^{\prime}$. For these experiments we take $\Gamma=1.5U_{0}$, $U_{0}^{\prime}=0.5U_{0}$. We then apply the result with the process $(S_{t},\Gamma+\sum_{k\leq t}(X_{k}-\widehat{\mu}_{k-1})(X_{k}-\widehat{\mu}_{k-1})^{\top}$ with the fixed matrix $U_{0}^{\prime}$ in place of $U_{0}$.
We compare Corollary 6.1 with the empirical Bernstein bound of Whitehouse et al. [63], which is given by (6.3). However, for the purposes of the experiments, we use the following version of the bound, stated by Whitehouse et al. [63] immediately after Theorem 4.1: For any $u>0$, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\leq 2\sqrt{L(V_{\tau})}+\frac{2L(V_{\tau})}{\gamma_{\min}(V_{\tau}+U_{0})},$ (C.1)
where
$L(V_{t})=\log\left(\frac{2}{\delta}\right)+\log\left(h\left(\log_{2}\left(\frac{\gamma_{\max}(V_{t}+U_{0})}{\gamma_{\min}(U_{0})}\right)\right)\right)+\log\left(2\sqrt{\kappa_{t}}\cdot N_{d-1}\left(\frac{1/4}{\sqrt{\kappa_{t}}}\right)\right),$
Despite being a relaxation of their main theorem for particular (unstated) hyperparameters, we find that it is actually tighter for the values of the hyperparameters that we tried, and more stable overall.